how to graph x=y^2-6y+8 This is parabola

Answer by lwsshak3(11628)

how to graph
x=y^2-6y+8
complete the square
x=(y^2-6y+9)+8-9
x=(y-3)^2-1
This is a parabola that opens rightwards
vertex: (-1,3)
x-intercept
set y=0
x=8
see graph below:
y=±(x+1)^.5+3